OpenStudy (anonymous):

y=ae^(bx) and passes through two points (-3,80) and 3,0.2) Determine the exact values of the sonstants a and b

OpenStudy (anonymous):

80 = ae^(-3b) 0.2 = ae^(0.2b) divide equation 1 by equation 2: 80/0.2 = ae^(-2b)/ae^(0.2b) so 400 = e^(-1.8b) so -1.8b = ln(400) so b = -ln(400)/1.8 thus a = 80*e^(3b) = 80*e^(-3ln(400)/1.8) = 80*e^(ln(400^(-3/1.8))) a = 80 * 400^(-3/1.8)

OpenStudy (anonymous):

not quite sure your numbers are correct. equation #2 should read 0.2 = ae^(3b)